# CAARMA: Class Augmentation with Adversarial Mixup Regularization

### v1

@@ Abstract
Speaker verification is a typical zero-shot learning task, where inference of unseen classes is performed by comparing embeddings of test instances to known examples. The models performing inference must hence naturally generate embeddings that cluster same-class instances compactly, while maintaining separation across classes. In order to learn to do so, they are typically trained on a large number of classes (speakers), often using specialized losses. However real-world speaker datasets often lack the class diversity needed to effectively learn this in a generalizable manner. We introduce CAARMA, a class augmentation framework that addresses this problem by generating synthetic classes through data mixing in the embedding space, expanding the number of training classes. To ensure the authenticity of the synthetic classes we adopt a novel adversarial refinement mechanism that minimizes categorical distinctions between synthetic and real classes. We evaluate CAARMA on multiple speaker verification tasks, as well as other representative zero-shot comparison-based speech analysis tasks and obtain consistent improvements: our framework demonstrates a significant improvement of 8% over all baseline models. Code for CAARMA will be released.
@@
Massa Baali, Xiang Li, Hao Chen, Rita Singh, Bhiksha Raj Carnegie Mellon University mbaali@cs.cmu.edu
@@ 1 Introduction
Popular approaches to address training data limitations often rely on data augmentation techniques. Methods such as AutoAugment Cubuk et al. (2019) and SpecAugment Park et al. (2019) generate new samples by modifying existing ones through transformations like geometric distortions, time warping, and frequency masking. However, while these techniques increase intra-class diversity, they do not introduce new classes, limiting their effectiveness in zero-shot learning scenarios. Of most relevance to our paper are mixup-based data augmentation techniques, e.g. Verma et al. (2019); Yun et al. (2019), that aim to enhance training by generating new samples through interpolation Han et al. (2021) of both the features and their labels. Regardless of the interpolation, yet these methods also do not generate new classes; they merely improve the generalization of the model by mapping mixed data to mixed-class labels. Still other methods use generative models such as VAEs, GANs, diffusion models etc. to generate authentically novel data to enhance the training Min et al. (2019); however these too are generally restricted to generating novel instances for known classes, limiting their effectiveness in zero-shot scenarios Pourpanah et al. (2022). Thus, while these approaches are generally very successful in improving generalization in classification problems, they fail at addressing the problem ZSL learning faces, that of increasing the number of classes themselves, leading to inconsistent generalization to unseen classes Xie et al. (2022).
@@ 2 Related-Work
Recent advancements in synthetic audio generation have emphasized the creation of diverse and high-quality datasets, crucial for training and evaluating audio-based AI models. A notable innovation is ConversaSynth, a framework utilizing large language models (LLMs) to generate synthetic conversational audio across varied persona settings (Gao et al., 2022). This method begins with generating text-based dialogues, which are then rendered into audio using text-to-speech (TTS) systems. The synthetic datasets produced are noted for their realism and topic variety, proving beneficial for tasks such as audio tagging, classification, and multi-speaker speech recognition. These capabilities make ConversaSynth a valuable tool for developing robust, adaptable AI models that can handle diverse audio data and complex conversational contexts. In the domain of speaker verification, SpeechMix introduces a novel method by mixing speech at the waveform level, carefully adjusting ratios to preserve the distinct characteristics of speaker identity (Jindal et al., 2020). However, like many other generative and augmentation techniques, SpeechMix primarily focuses on manipulating known speaker voices rather than generating new identities, thereby limiting its utility for enhancing speaker diversity critical for effective zero-shot learning. Synthio employs a unique approach by using text-to-audio (T2A) diffusion models to augment small-scale audio classification datasets (Joassin and Alvarez-Melis, 2022). It enhances compositional diversity and maintains acoustic consistency by aligning T2A-generated synthetic samples with the original dataset using preference optimization. Additionally, exploring style transfer in synthetic audio, recent work by Ueda et al. employs a VITS-based voice conversion model, conditioned on the fundamental frequency (F0), to produce expressive variations from neutral speaker voices (Ueda et al., 2024). This method achieves cross-speaker style transfer in a FastPitch-based TTS system, incorporating a style encoder pre-trained on timbre-perturbed data to prevent speaker leakage. This technique enhances the utility of synthetic data in applications requiring rich stylistic diversity. These developments underline the increasing sophistication of synthetic audio generation techniques, from multi-speaker conversations in ConversaSynth to Synthio’s optimized T2A augmentation for classification, and cross-speaker style transfers with VITS-based models. They collectively demonstrate the power of synthetic data to enrich audio model training by adding diversity and realism. However, these methods still face limitations in generating entirely new speaker identities, which is critical for expanding the range of recognizable voices in speaker verification systems. CAARMA addresses this gap by directly mixing in the embedding space, creating synthetic speakers that enhance zero-shot learning capabilities. CAARMA not only preserves speaker characteristics but also significantly expands the diversity of speaker identities, offering a superior solution for training more robust and adaptable speaker verification systems.
@@ 3.1 Overview
As mentioned in Section 1, ZSL models learn their ability to compactly cluster same-class embeddings while maintaining separation between classes primarily through exposure to a large number of classes during training; the more classes they are exposed to in training, the better they are able to generalize to unseen classes. In the speaker verification setting, this translates to training the model with recordings from a large number of speakers; the more the number of training speakers the better the model generalizes. To improve this generalization CAARMA attempts to increase the number of speakers by creating synthetic speakers while training. Synthetic speakers may be created through generative methods such as Cornell et al. (2024); however this approach does not scale. Instead, CAARMA creates them through a simple mixup strategy, as convex combinations of real speakers, with a key distinction: the mixup is performed in the embedding space, where the classes are expected to form compact (and generally convex) clusters. In order to ensure that these synthetic speakers are indeed representative of actual speakers, CAARMA utilizes a semantic discriminator, a discriminator which attempts to distinguish between synthetic and real speakers: if this discriminator is fooled, the synthetic speakers are statistically indistinguishable from real ones.
@@ 3.2 Framework
Our framework consists of three main components: an encoder for embedding generation, a synthetic label mixup mechanism for class augmentation, and an adversarial training scheme with a semantic discriminator. Figure 2 illustrates the complete pipeline of our approach. The process begins with a waveform input that is transformed into a Mel-spectrogram. This spectrogram serves as input to the encoder $\mathcal{E}$, which generates embeddings $e$ that capture discriminative speaker characteristics. These embeddings undergo our SL-Mixup strategy, which generates synthetic embeddings $e_{\text{syn}}$ by mixing embeddings $e$ based on their closest neighbor weights $W$. Each synthetic embedding receives a corresponding synthetic label $ID_{\text{syn}}$ within the mini-batch. The framework employs two primary loss functions: the encoder loss $\mathcal{L}_{\text{real}}$ for original embeddings and the synthetic loss $\mathcal{L}_{\text{syn}}$ for synthetic embeddings. A Self-Supervised Learning (SSL) model serves as the semantic discriminator to distinguish between real ($R$) and synthetic ($S$) embeddings. A discriminator loss $\mathcal{L}_{D}$ guides the discriminator to maximally distinguish between $R$ and $S$. On the other hand, a generator loss $\mathcal{L}_{\text{gen}}$ guides the encoder to “fool” the discriminator, so that it perceives no distinction between real and synthetic embeddings.
@@ 3.5 Adversarial Training
3.6 Semantic Discriminator
To enhance the discriminative power of our framework, we incorporate a self-supervised model (HuBERT) Hsu et al. (2021) as a semantic discriminator. This discriminator leverages the pre-trained representations to provide richer gradients during adversarial training, improving the stability and quality of the learned embeddings. The semantic discriminator processes embeddings through an adapter module that projects them into a compatible feature space. The adapter consists of a down-projection layer with spectral normalization, followed by fully connected layers with GELU activation Hendrycks and Gimpel (2016). We employ skip connections and layer normalization to ensure stable training. The discriminator extracts features from multiple HuBERT Hsu et al. (2021) layers (7, 9, 11, and 12) to capture diverse speaker characteristics. These features are combined using learnable weights and processed through a residual classification block with spectral normalization and LeakyReLU activation Xu (2015) to determine whether an embedding is real or synthetic.
@@ 4.1 Datasets
  • Emotion Recognition: We use the DailyTalk dataset to classify the emotional state of a speaker based on speech utterances. This dataset contains speech samples labeled with seven distinct emotion categories. To maintain consistency with other datasets, we resample all recordings to 16 kHz before processing.
@@
ID Dataset Classes Utterances
1 VoxCeleb1 1211 1,087135
2 VoxCeleb2 5994 153,516
3 HowFarAreyou 3 30,00
4 DailyTalk 7 16,600
Table 2: EER Results for MFA Conformer baseline, Adversarial Training (AT), Semantic Discriminator (SD), and Synthetic Loss $L_{syn}$ using VoxCeleb1 for the SV tasks.
ID $L_{syn}$ AT SD Results
@@ 4 DailyTalk 7 16,600
Encoder Baseline AT AT+$\mathcal{L}_{syn}$
@@
Table 4: Ablation study of different hidden layers for Semantic Discriminator (SD) reporting EER (%) and minDCF.
@@
Model # Parameters EER(%) minDCF
ResNet34 23.2M 1.03 0.112
ECAPA-TDNN 20.8M 0.82 0.112
MFA-Conformer 19.8M 0.86 0.066
MFA-Adversarial 19.8M 0.81 0.036
@@ 5.1 Speaker Verification Task
We evaluate our models on VoxCeleb1-O, the official test set of VoxCeleb1.
@@ 5.1.1 Small Scale
Using the MFA Conformer as the Encoder, we conduct several experiments (Table 2) reporting the Equal Error Rate (EER). The addition of synthetic loss alone (Model $ID_{2}$) yield slight improvements over the baseline (Model $ID_{1}$). More substantial gains were achieved through adversarial training (Model $ID_{3}$), with further improvements observed when incorporating the semantic discriminator (Model $ID_{4}$). The best performance was achieved by Model $ID_{6}$, which combined all three components: synthetic loss, semantic discriminator, and adversarial training.
To further validate the generalizability of our approach, we implement it with an alternative speaker encoder. As shown in Table 3, the combination of adversarial training and semantic discriminator improved performance by 6.56% compared to the baseline. Adding synthetic loss further enhanced the improvement to 8.29%.
We conduct an ablation study to identify the most informative hidden layers for speaker representation. Table 4 presents the EER and minimum Detection Cost Function (minDCF) across various layer configurations. Our analysis revealed that layers 7, 9, 11, and 12 provide the most effective speaker characteristics representation, suggesting that later layers capture more valuable speaker-specific information.
@@ 5.1.2 Large Scale
To demonstrate the scalability of our framework, we conduct experiments on models trained on both VoxCeleb1 and VoxCeleb2, as a large-scale dataset. As shown in Table 5, we incorporate the MFA Conformer as the encoder within our adversarial training framework, leveraging its strong baseline performance on smaller datasets. The results indicate that our adversarial refinement approach scales effectively to larger datasets, consistently improving both EER and minDCF metrics.

### v3

@@ Abstract
Speaker verification is a typical zero-shot learning task, where inference of unseen classes is performed by comparing embeddings of test instances to known examples. The models performing inference must hence naturally generate embeddings that cluster same-class instances compactly, while maintaining separation across classes. In order to learn to do so, they are typically trained on a large number of classes (speakers), often using specialized losses. However real-world speaker datasets often lack the class diversity needed to effectively learn this in a generalizable manner. We introduce CAARMA, a class augmentation framework that addresses this problem by generating synthetic classes through data mixing in the embedding space, expanding the number of training classes. To ensure the authenticity of the synthetic classes we adopt a novel adversarial refinement mechanism that minimizes categorical distinctions between synthetic and real classes. We evaluate CAARMA on multiple speaker verification tasks, as well as other representative zero-shot comparison-based speech analysis tasks and obtain consistent improvements: our framework demonstrates a significant improvement of 8% over all baseline models. The code is available at: https://github.com/massabaali7/CAARMA/
@@
Massa Baali, Xiang Li, Hao Chen, Syed Abdul Hannan, Rita Singh, Bhiksha Raj Carnegie Mellon University mbaali@cs.cmu.edu
@@ 1 Introduction
Popular approaches to address training data limitations often rely on data augmentation techniques. Methods such as AutoAugment Cubuk et al. (2019) and SpecAugment Park et al. (2019) generate new samples by modifying existing ones through transformations like geometric distortions, time warping, and frequency masking. However, while these techniques increase intra-class diversity, they do not introduce new classes, limiting their effectiveness in zero-shot learning scenarios. Of most relevance to our paper are mixup-based data augmentation techniques, e.g. Verma et al. (2019); Yun et al. (2019), that aim to enhance training by generating new samples through interpolation Han et al. (2021) of both the features and their labels. Regardless of the interpolation, yet these methods also do not generate new classes; they merely improve the generalization of the model by mapping mixed data to mixed-class labels. Still other methods use generative models such as Variational Autoencoders (VAEs), Generative Adversarial Networks (GANs), diffusion models etc. to generate authentically novel data to enhance the training Min et al. (2019); however these too are generally restricted to generating novel instances for known classes, limiting their effectiveness in zero-shot scenarios Pourpanah et al. (2022). Thus, while these approaches are generally very successful in improving generalization in classification problems, they fail at addressing the problem ZSL learning faces, that of increasing the number of classes themselves, leading to inconsistent generalization to unseen classes Xie et al. (2022).
@@ 2 Related-Work
Recent advancements in synthetic audio generation have emphasized the creation of diverse and high-quality datasets, crucial for training and evaluating audio-based AI models. A notable innovation is ConversaSynth, a framework utilizing large language models (LLMs) to generate synthetic conversational audio across varied persona settings (Gao et al., 2022). This method begins with generating text-based dialogues, which are then rendered into audio using text-to-speech (TTS) systems. The synthetic datasets produced are noted for their realism and topic variety, proving beneficial for tasks such as audio tagging, classification, and multi-speaker speech recognition. These capabilities make ConversaSynth a valuable tool for developing robust, adaptable AI models that can handle diverse audio data and complex conversational contexts. In the domain of speaker verification, SpeechMix introduces a novel method by mixing speech at the waveform level, carefully adjusting ratios to preserve the distinct characteristics of speaker identity (Jindal et al., 2020). However, like many other generative and augmentation techniques, SpeechMix primarily focuses on manipulating known speaker voices rather than generating new identities, thereby limiting its utility for enhancing speaker diversity critical for effective zero-shot learning. In the context of speaker verification, discriminative neural clustering Li et al. (2021) employs clustering techniques to enhance speaker diarization by learning discriminative embeddings, but it does not address class diversity through synthetic class generation as CAARMA does. Synthio employs a unique approach by using text-to-audio (T2A) diffusion models to augment small-scale audio classification datasets (Joassin and Alvarez-Melis, 2022). It enhances compositional diversity and maintains acoustic consistency by aligning T2A-generated synthetic samples with the original dataset using preference optimization. Additionally, exploring style transfer in synthetic audio, recent work by Ueda et al. employs a VITS-based voice conversion model, conditioned on the fundamental frequency (F0), to produce expressive variations from neutral speaker voices (Ueda et al., 2024). This method achieves cross-speaker style transfer in a FastPitch-based TTS system, incorporating a style encoder pre-trained on timbre-perturbed data to prevent speaker leakage. This technique enhances the utility of synthetic data in applications requiring rich stylistic diversity. These developments underline the increasing sophistication of synthetic audio generation techniques, from multi-speaker conversations in ConversaSynth to Synthio’s optimized T2A augmentation for classification, and cross-speaker style transfers with VITS-based models. They collectively demonstrate the power of synthetic data to enrich audio model training by adding diversity and realism. However, these methods still face limitations in generating entirely new speaker identities, which is critical for expanding the range of recognizable voices in speaker verification systems. CAARMA addresses this gap by directly mixing in the embedding space, creating synthetic speakers that enhance zero-shot learning capabilities. CAARMA not only preserves speaker characteristics but also significantly expands the diversity of speaker identities, offering a superior solution for training more robust and adaptable speaker verification systems.
@@ 3.1 Overview
As mentioned in Section 1, ZSL models learn their ability to compactly cluster same-class embeddings while maintaining separation between classes primarily through exposure to a large number of classes during training; the more classes they are exposed to in training, the better they are able to generalize to unseen classes. In the speaker verification setting, this translates to training the model with recordings from a large number of speakers; the more the number of training speakers the better the model generalizes. To improve this generalization CAARMA attempts to increase the number of speakers by creating synthetic speakers while training. Synthetic speakers may be created through generative methods such as Cornell et al. (2024); however this approach does not scale. Instead, CAARMA creates them through a simple mixup strategy, as convex combinations of real speakers, with a key distinction: the mixup is performed in the embedding space, where the classes are expected to form compact (and generally convex) clusters. In order to ensure that these synthetic speakers are indeed representative of actual speakers, CAARMA utilizes a mixup discriminator, a discriminator which attempts to distinguish between synthetic and real speakers: if this discriminator is fooled, the synthetic speakers are statistically indistinguishable from real ones.
@@ 3.2 Framework
Our framework consists of three main components: an encoder for embedding generation, a synthetic label mixup mechanism for class augmentation, and an adversarial training scheme with a semantic discriminator. Figure 2 illustrates the complete pipeline of our approach. The process begins with a waveform input that is transformed into a Mel-spectrogram. This spectrogram serves as input to the encoder $\mathcal{E}$, which generates embeddings $e$ that capture discriminative speaker characteristics. These embeddings undergo our SL-Mixup strategy, which generates synthetic embeddings $e_{\text{syn}}$ by mixing embeddings $e$ based on their closest neighbor weights $W$. Each synthetic embedding receives a corresponding synthetic label $ID_{\text{syn}}$ within the mini-batch. The framework employs two primary loss functions: the encoder loss $\mathcal{L}_{\text{real}}$ for original embeddings and the synthetic loss $\mathcal{L}_{\text{syn}}$ for synthetic embeddings. A Self-Supervised Learning (SSL) model serves as the mixup discriminator to distinguish between real ($R$) and synthetic ($S$) embeddings. A discriminator loss $\mathcal{L}_{D}$ guides the discriminator to maximally distinguish between $R$ and $S$. On the other hand, a generator loss $\mathcal{L}_{\text{gen}}$ guides the encoder to “fool” the discriminator, so that it perceives no distinction between real and synthetic embeddings.
@@ 3.5 Adversarial Training
3.6 Mixup Discriminator
To enhance the discriminative power of our framework, we incorporate a self-supervised model (HuBERT) Hsu et al. (2021) as a mixup discriminator. This discriminator leverages the pre-trained representations to provide richer gradients during adversarial training, improving the stability and quality of the learned embeddings. The semantic discriminator processes embeddings through an adapter module that projects them into a compatible feature space. The adapter consists of a down-projection layer with spectral normalization, followed by fully connected layers with GELU activation Hendrycks and Gimpel (2016). We employ skip connections and layer normalization to ensure stable training. The discriminator extracts features from multiple HuBERT Hsu et al. (2021) layers (7, 9, 11, and 12) to capture diverse speaker characteristics. These features are combined using learnable weights and processed through a residual classification block with spectral normalization and LeakyReLU activation Xu (2015) to determine whether an embedding is real or synthetic.
@@ 4.1 Datasets
  • Emotion Recognition: We use the DailyTalk dataset to classify the emotional state (anger, disgust, fear, happiness, sadness, surprise, neutral) of a speaker based on speech utterances. This dataset contains speech samples labeled with seven distinct emotion categories. To maintain consistency with other datasets, we resample all recordings to 16 kHz before processing.
@@
ID Dataset Classes Utterances
1 VoxCeleb1 1211 153,516
2 VoxCeleb2 5994 1,087135
3 HowFarAreyou 3 3,000
4 DailyTalk 7 16,600
@@
ID $L_{syn}$ AT MD Results
@@
Table 2: EER Results for MFA Conformer baseline, Adversarial Training (AT), Mixup Discriminator (MD), and Synthetic Loss $L_{syn}$ using VoxCeleb1 for the SV tasks.
[…]
@@
Table 4: Ablation study of different hidden layers for Mixup Discriminator (MD) reporting EER (%) and minDCF.
ID Model # Parameters EER(%) minDCF Utterances # Speakers
1 MFA-Conformer 19.8M 12.82 0.770 12,335 100
2 MFA-Adversarial 19.8M 11.83 0.790 12,335 100
3 MFA-Conformer 19.8M 0.86 0.066 1,240,651 7,205
4 MFA-Adversarial 19.8M 0.81 0.036 1,240,651 7,205
@@ 4.2.3 Implementation Details
Computational Complexity:
CAARMA introduces no additional computational cost during inference, as the discriminator is not used, and inference remains identical to the baseline model. During training, the primary computational overhead arises from two components: (1) the forward and backward passes through the discriminator, and (2) the computation of the AM-Softmax loss for both real and synthetic (mixup) data. Since synthetic embeddings are dynamically generated from real-data embeddings, no additional embedding computations are required. In our experiments, synthetic data is generated in a 1:1 ratio with real data, effectively quadrupling the compute time for the loss calculations (doubling for AM-Softmax on real and synthetic data, and doubling for the discriminator). However, as the loss computation constitutes a minor fraction of the overall training cost, the total increase in training time remains modest. Additionally, since mixup data is computed dynamically, the memory overhead is negligible.
@@ 5.1 Speaker Verification Task
We evaluate our models on VoxCeleb1-O, the official test set of VoxCeleb1. For evaluating the performance, we use the Equal Error Rate (EER) and minimum Detection Cost Function (minDCF). EER represents the point where the false acceptance rate equals the false rejection rate, providing a single measure of verification accuracy (lower is better). The minDCF quantifies the cost of detection errors, balancing false positives and false negatives, with lower values indicating better performance. These metrics assess the model’s ability to distinguish between same-speaker and different-speaker pairs effectively.
@@ 5.1.1 Small Scale
Using the MFA Conformer as the Encoder, we conduct several experiments (Table 2) reporting the EER. The addition of synthetic loss alone (Model $ID_{2}$) yield slight improvements over the baseline (Model $ID_{1}$). More substantial gains were achieved through adversarial training (Model $ID_{3}$), with further improvements observed when incorporating the mixup discriminator (Model $ID_{4}$). The best performance was achieved by Model $ID_{6}$, which combined all three components: synthetic loss, mixup discriminator, and adversarial training.
To further validate the generalizability of our approach, we implement it with an alternative speaker encoder. As shown in Table 3, the combination of adversarial training and mixup discriminator improved performance by 6.56% compared to the baseline. Adding synthetic loss further enhanced the improvement to 8.29%.
We conduct an ablation study to identify the most informative hidden layers for speaker representation. Table 4 presents the EER and minDCF across various layer configurations. Our analysis revealed that layers 7, 9, 11, and 12 provide the most effective speaker characteristics representation, suggesting that later layers capture more valuable speaker-specific information.
@@ 5.1.2 Large Scale
To demonstrate scalability, we train on the combined VoxCeleb1 and VoxCeleb2 datasets, creating a substantially larger training corpus. As summarized in Table 5, the MFA-Conformer encoder (Model $ID_{3}$) achieves strong baseline performance, which is further enhanced through adversarial refinement (Model $ID_{4}$).
Specifically, adversarial training reduces the EER from 0.86% to 0.81% and nearly halves the minDCF from 0.066 to 0.036. These improvements are consistent with the small-scale trends, confirming that our framework scales robustly to larger and more diverse speaker populations.
Importantly, we note that even in the limited-diversity setting of VoxCeleb1—where only 100 speakers are present—our approach still yields consistent gains (Model IDs 1–2). This suggests that the framework is not merely leveraging broader class coverage in large-scale datasets, but is also effective in scenarios with constrained speaker diversity. In practice, this means that our method improves speaker verification robustness both under resource-rich conditions and under more restrictive data availability.
Overall, these results highlight that adversarial training is beneficial across training regimes: it generalizes well to large-scale, diverse datasets while still offering tangible improvements when diversity is limited.